# TIAM - A Metric for Evaluating Alignment in Text-to-Image Generation

### v1

@@ 4.1 Preliminary: performance drop with 2 objects
We report a first experiment to illustrate the general framework of our study.
We consider various labels from COCO (the exhaustive list is in supplementary material B), leading to a set of $|\mathcal{O}|=24$ object labels. For all possible couples of labels $(o_{i},o_{j})$ with $o_{i}\neq o_{j}$, we make the prompt “a photo of $o_{i}$ and $o_{j}$” (managing the determinant as expected) and generate 64 synthetic images by changing the random seed $\chi$ and estimate the alignment of the prompt with each image with TIAM. As a reference, we also conduct the same experiment with 64 images generated from the simpler prompt “a photo of $o_{i}$”.
We report the result in Tab 1 for all the models considered. The models consistently succeed in generating images with a single object, with a score above 0.95. However, they struggle to generate simultaneously two objects with at most 66% of images correctly generated, while the prompts are minimally simple. In the recent literature in the field, the T2I models usually exhibit a good quantitative score of rendering, e.g. in terms of inception score or FID, which our experience does not call into question. In the vein of Hinz et al. [10] this experiment only shows that there is a problem with alignment between the content of the generated image and the prompt. More precisely than [10], our experiment specifically quantifies to which extent this drop in alignment performance is due to the presence of multiple concepts in the prompt. For the prompts with two objects, we report in Tab 2 the number of times the objects are actually accurately generated with regard to their position in the prompt (the score in Tab 1 requires the presence of both objects to be valid). We observe a tendency for the first $o_{1}$ to be more prevalent than the second $o_{2}$ across all models, a phenomenon that is further investigated in section 4.3.
By comparing the performances of the models, we notice that A&E inference globally improves the score with a large margin for SD v1.4. From Tab 2, it seems that the overperformance of A&E is due to its ability to enhance the occurrence rate of $o_{2}$ by enforcing the minimum excitation of cross-attention maps.
@@ 4.2 Importance of the random seed
We consider the same 24 labels as in subsection 4.1 and the prompt “a photo of $o_{i}$ and $o_{j}$”. We also generate 64 synthetic images by changing the random seed $\chi$ and compute $f\left(G(\chi,t(z)),y(z)\right)$ for each prompt and corresponding image. However, this time, the 64 seeds are the same for all the $(o_{i},o_{j})$ [^5] Actually, it was the case in subsection 4.1 for the sake of comparison and coherency and we aggregate the performance per seed, for all possible prompts and images, leading to $24\times 23=552$ estimations of text-image alignment per seed.
@@
The results in Fig 3 show that, for all 6 models, the performance varies a lot with regard to the random seed. These results, which to the best of our knowledge have never been identified in the literature of the T2I generative models, may seem surprising. With such models, one usually expects that all seeds have the same chance to generate the specified elements. Actually, the recent tendency rather consists of engineering finely the prompt [18] to optimize the output of T2I models. On the opposite of this tendency, our method opens up new perspectives toward a complementary optimization based on the choice of “performing seeds” (see subsection 4.5).
@@
Overall, there exists a significant disparity between the best and worst seeds. When considering the interquartile ranges and the min-max range through the box-plots of Fig 3, it appears that the difference in average performance between models is much less significant than their own inner difference due to the variation of the seeds.
@@
Obviously, enhancing a dependence “to the seed” is convenient from a practical point of view, but wrong in all strictness since the seeds are drawn from a Gaussian noise in inference. Being more rigorous requires reminding the diffusion models training process [11]. It consists in learning the reverse process of a fixed Markov chain of length $T$ with models that can be interpreted as an equally weighted sequence of denoising autoencoders. These last are trained to denoise their input, considered as a noisy version of an input training image $I$. In the case of latent diffusion models [24], the process is embedded into the latent space of a VAE (encoder $\mathcal{E}$ + decoder $\mathcal{D}$), such that the autoencoders are U-net networks $\epsilon_{\theta}(x_{t},t)$, $t=1\dots T$ that denoise the latent code $x_{t}$ by minimizing the loss $\mathbb{E}_{\mathcal{E}(I),\epsilon\sim\mathcal{N}(0,1)}\left[||\epsilon-\epsilon_{\theta}(x_{t},t)||_{2}^{2}\right]$. During inference, the process starts from a latent code $\chi_{T}\sim\mathcal{N}(0,1)$, denoises it with the U-net $\epsilon_{\theta}$ to get the final latent code $\chi_{0}$, then obtains the synthetic image with VAE decoder as $\mathcal{D}(\chi_{0})$. This last is thus not just dependent on the random seed $\chi_{T}$ but on the U-net parameter $\theta$ as well. These parameters are learned to denoise the latent codes at any step of the diffusion since the reparametrization trick [11] allows to express the final code directly as:
@@
where $\epsilon\sim\mathcal{N}(0,1)$ and the $\alpha_{t}$ depends on the variance of the diffusion process noise (see [11] and the appendix B of [24] for details). Ideally, $\bar{\alpha_{T}}$ is close to one, such that a synthetic image should result from an almost perfect Gaussian noise. However, in practice the process is not perfect, thus the final $\chi_{0}$ obtained for a given $\chi_{T}\sim\mathcal{N}(0,1)$ at inference time is only an approximation of the “ideal $\chi_{0}$” that one could expect with a perfect optimization [6]. Moreover, since each model is trained independently, its latent space is structured in the same vein and the “path” of the $\chi_{t}$ in the latent space may strongly differ from one model to another. Hence, starting from the same $\chi_{T}\sim\mathcal{N}(0,1)$, they result in quite different $\chi_{0}$ while trying to tend to an ideal one. As a consequence, the “good” and “bad” seeds are specific to each model.
Figure 3: TIAM aggregate per seed for 64 seeds. We show that some started latent noise tends to not be converted to an image with two entities regardless of the entities. “+” shows the mean.
The results are contingent upon the seed, indicating that based on the prompt’s specifications, it appears possible to identify initial noises that are more likely to exhibit multiple objects. It highlights the ongoing need for further advancements in reducing reliance on latent variables. To obtain a robust score and avoid the possible impact of the high variability of the seed’s success rate, we define a minimum of images to generate per prompt. We establish it is adequate to generate 32 images per prompt to obtain a robust score. The corresponding experiment is reported in the supplementary material E.
@@ 4.3 Catastrophic Neglect
This section examines the behavior of the models as the number of object sets increases and the role of the order of objects in the prompt.
@@
When examining the occurrence of different objects, similar trends can be observed as in the preliminary experiment (Tab 2). Specifically, the initial objects in the template tend to appear more frequently than objects inserted subsequently. The results for the template containing four objects are presented in Fig 5 (results for two and three objects in supplementary material D). This reinforces the observation that the concept that is expressed earlier in the prompt has more chances to appear in the final image.
For SD, conditioned by CLIP text-encoders, we hypothesize that the decreasing trend in the occurrence of objects as their position becomes more distant in the prompt is partly due to the auto-regressive nature of the encoder. During self-attention, tokens only receive context from the elements to their left (beginning of the prompt). As a result, tokens are devoid of the contexts of subsequent objects, while the following objects carry the context of the earlier words. This leads to a more ambiguous signal from distant tokens for U-Net models.
The explanation is nevertheless less clear with the T5 encoder, where all words have access to each other during self-attention. Used in IF, one can see that the third and fourth words have the same chance to appear, but it is significantly less than the object in the second position, itself below the first concept expressed in the prompt.
We hypothesize that during cross-attention, the models have learned to give more importance to tokens with earlier positions due to the training data, which typically places important elements related to the image at the beginning of the caption. Asserting this explanation would nevertheless require significant work to analyze the original training dataset used to pre-trained the six generative models, which is out of the scope of this paper.
@@
Figure 5: The proportion of occurrences of each object, based on its positions in the prompt. The template of the prompt count 4 objects.
@@
Finally, we investigated the impact of semantic relationships between objects within the prompt on the ability of the model to generate both of them. Considering a template with two objects, our hypothesis was that the T2I model would fail more often to represent on the same image two objects that are semantically linked. We considered 28 COCO labels from three macro-classes (vehicles, animals, and foods) and generate images using a template with 2 objects. From the resulting TIAM scores, we derived a dissimilarity metric between all objects (see supplementary material H) and project all the labels with Multidimensional Scaling (MDS). The resulting projections (Fig. 18 for SD 1.4) can be interpreted such that the closer two labels are, the more challenging it becomes for the model to represent them together. Across all models, we observe labels from the same class being clustered together, particularly for animals and vehicles. This suggests the presence of semantic proximity. We measured the semantic distance between both named objects with various methods (such as Wu-Palmer, Cosine similarity of CLIP text embeddings or even attention’s key-value representation) for SD 1.4 and SD 2 but the correlation with the TIAM score was only slightly negative. This indicates that the semantic linking has probably either a small or indirect link to the alignment performance, but in any case, further researches need to be conducted to clarify this point.
@@ 4.4 Attribute binding
We start by computing TIAM regardless of the correctness of the attribute and report the results in Tab 3. Without attributes (✗) the scores are close to those of Tab 1 and adding attributes (✓) has a limited impact when one object only is requested. However, it significantly impacts the performance when two objects are present in the prompt.
@@
Figure 8: TIAM per color and per object.
@@
When TIAM is estimated by taking into account both the object and color ground truth, we observe that the models fail to assign the appropriate colors to the objects, even when these last are present in the image (Fig. 7). For instance, in the case of SD 1.4, objects are detected correctly in 95% of single-object prompts, but only approximately 45% of them have the requested colors.
In Fig. 8, we report the attribution scores per color and per object to analyze deeper the attributing ability of the models. We see that models fail to generalize the binding to objects when it is uncommon to observe them in that particular color. As expected, it is indeed easier to assign colors to cars and refrigerators compared to animals. The models have likely been trained on photos of cars and refrigerators in various colors during training as it is more common to come across a green car than a green giraffe.
Following the same approach as in section 4.3, we also analyze the results in relation to the position $i$ of attribute $a_{i}$ and object $o_{i}$ in the prompt involving 2 objects. In that case, the models succeed by a large margin to generate and bind the first object. However, knowing that the first object is more often generated we compute a binding success rate, which is the score of correctly attributed objects among the correctly generated objects (Fig. 9). The $o_{2}$ objects, however, continue to be less attributed. This reinforces the finding that the first object in the prompt has a greater influence on the final generation.
@@
In the supplementary material I, we report the results of the binding success rate differentiated by colors for attributes in the first position and attributes in the second position. We observed that the models face greater difficulty in assigning green and blue colors when two objects are involved (parallel with a single label case). It is worth noting that IF performs better than other models.
@@ 4.5 Toward noise mining?
To highlight the importance of noise performance, we selected seeds based on their TIAM score. In Fig. 10 we present qualitative results using some of the worst and best seeds found with our approach, and prompts with objects and attributes that were not considered in our study.
Both with SD 2 and IF, the two objects are better represented in the image resulting from the “good” seed, showing that our method has the ability to find seeds that generalize to objects out of the domain of objects used to determine the best seeds.
@@
We conducted the same seed selection process for both our worst and best seeds using colors in the prompt (Fig. 11). Once again the images resulting from the best seeds we identified with TIAM better reflect the prompt. With the “bad” seed, SD 2 suffers both from attribute leaking or binding (the “blue moon” is yellow, and the “red lion” is blue), and for both models, the images resulting from the “good” seed are more aligned. This emphasizes the dependency on started latent noise present at the beginning of the prompt, which remains a prominent factor of performance.
@@
5 Conclusion and Perspective
We proposed a new metric to automatically quantify the performances of T2I models in terms of prompt-image alignment. Contrary to previous research efforts, it is based on prompt templates, that allow a finer analysis with regard to the syntax of the prompt. Hence we showed that the alignment performances of most T2I models drop significantly with the number of objects specified in the prompt and that effect is even more critical for color attribution. Our metric also allows us to study the influence of the input seed at inference. We showed that there exist some seeds that systematically result in better output images than others, and that it generalizes on objects out of the set used to determine them. It draws possible future research toward the mining of such “good seeds”, as a complementary activity to prompt engineering to optimize the outputs of T2I models.

### v2

@@ 4.1 Preliminary: performance drop with 2 objects
We report a first experiment to illustrate the general framework of our study. We consider various labels from COCO (the exhaustive list is in Supp. Mat. B), leading to a set of $|\mathcal{O}|=24$ object labels. For all possible pairs of labels $(o_{i},o_{j})$ with $o_{i}\neq o_{j}$, we make the prompt “a photo of $o_{i}$ and $o_{j}$” (managing the determinant as expected) and generate 64 synthetic images by changing the random seed $\chi$ and estimate the alignment of the prompt with each image using TIAM. As a reference, we also conduct the same experiment with 64 images generated from the simpler prompt “a photo of $o_{i}$”.
We report the results in Tab. 1 for all the models considered. The models consistently succeed in generating images with a single object, with a score above 0.95. However, they struggle to generate simultaneously two objects, with at most 66% of images correctly generated, while the prompts are minimally simple. In the recent literature, the T2I models usually exhibit a good quantitative rendering score, e.g. in terms of inception score or FID, which our experience does not call into question. In the vein of Hinz et al. [19], this experiment only shows that there is a problem with alignment between the content of the generated image and the prompt. More precisely than [19], our experiment specifically quantifies to which extent this drop in alignment performance is due to the presence of multiple concepts in the prompt. For the prompts with two objects, we report in Tab. 2 the number of times the objects are actually accurately generated with regard to their position in the prompt (the score in Tab. 1 requires the presence of both objects to be valid). We observe a tendency for the first $o_{1}$ to be more prevalent than the second $o_{2}$ across all models, a phenomenon that is further investigated in Section 4.3.
We notice that A&E inference globally improves the score with a large margin for SD v1.4. From Tab. 2, it seems that the overperformance of A&E is due to its ability to enhance the occurrence rate of $o_{2}$ by enforcing the minimum excitation of cross-attention maps.
@@ 4.2 Importance of the random seed
We consider the same 24 labels as in Section 4.1 and the prompt “a photo of $o_{i}$ and $o_{j}$”. We also generate 64 synthetic images by changing the random seed $\chi$ and compute $f\left(G(\chi,t(z)),y(z)\right)$ for each prompt and the corresponding image. However, this time, the 64 seeds are the same for all the $(o_{i},o_{j})$ [^2] Actually, it was the case in Section 4.1 for the sake of comparison and coherency. and we aggregate the performance per seed, for all possible prompts and images, leading to $24\times 23=552$ estimations of text-image alignment per seed.
@@
The results in Fig. 3 show that, for all 6 models, the performance varies a lot with regard to the random seed. These results, which to the best of our knowledge have never been identified in the literature of the T2I generative models, may seem surprising. With such models, one usually expects that all seeds have the same chance to generate the specified elements. Actually, the recent tendency rather consists of engineering finely the prompt [34] to optimize the output of T2I models. On the opposite of this tendency, our method opens up new perspectives toward a complementary optimization based on the choice of “performing seeds” (see Section 4.6).
@@
Overall, there exists a significant disparity between the best and worst seeds. When considering the interquartile ranges and the min-max range through the box-plots of Fig. 3, it appears that the difference in average performance between models is much less significant than their own inner difference due to the variation of the seeds.
@@
Obviously, enhancing a dependence “to the seed” is convenient from a practical point of view but wrong in all strictness since the initial noises are drawn from a Gaussian at inference. Being more rigorous requires reminding the diffusion models training process [20]. It consists of learning the reverse process of a fixed Markov chain of length $T$ with models that can be interpreted as an equally weighted sequence of denoising autoencoders. The latters are trained to denoise their input, considered as a noisy version of an input training image $I$. In the case of latent diffusion models [40], the process is embedded into the latent space of a VAE (encoder $\mathcal{E}$ + decoder $\mathcal{D}$), such that the autoencoders are U-net networks $\epsilon_{\theta}(x_{t},t)$, $t=1\dots T$ that denoise the latent code $x_{t}$ by minimizing the loss $\mathbb{E}_{\mathcal{E}(I),\epsilon\sim\mathcal{N}(0,1)}\left[||\epsilon-\epsilon_{\theta}(x_{t},t)||_{2}^{2}\right]$. During inference, the process starts from a latent code $\chi_{T}\sim\mathcal{N}(0,1)$, denoises it with the U-net $\epsilon_{\theta}$ to get the final latent code $\chi_{0}$, then obtains the synthetic image with VAE decoder as $\mathcal{D}(\chi_{0})$. This last is thus not just dependent on the random seed $\chi_{T}$ but on the U-net parameter $\theta$ as well. These parameters are learned to denoise the latent codes at any step of the diffusion since the reparametrization trick [20] allows to express the final code directly as:
@@
where $\epsilon\sim\mathcal{N}(0,1)$ and the $\alpha_{t}$ depends on the variance of the diffusion process noise (see [20] and appendix B of [40] for details). Ideally, $\bar{\alpha_{T}}$ is close to one, such that a synthetic image should result from an almost perfect Gaussian noise. However, in practice the process is not perfect; thus the final $\chi_{0}$ obtained for a given $\chi_{T}\sim\mathcal{N}(0,1)$ at inference time is only an approximation of the “ideal $\chi_{0}$” that could be expected with a perfect optimization [9]. Moreover, since each model is trained independently, its latent space is structured in the same vein and the “path” of the $\chi_{t}$ in the latent space may strongly differ from one model to another. Hence, starting from the same $\chi_{T}\sim\mathcal{N}(0,1)$, they result in quite different $\chi_{0}$ while trying to tend to an ideal one. As a consequence, the “good” and “bad” seeds are specific to each model.
Figure 3: TIAM aggregated per seed for 64 seeds. We show that some starting noise tends to not be converted to an image with two entities regardless of the entities. “+” shows the mean.
The results depend on the seed, indicating that based on the prompt’s specifications, it appears possible to identify initial noises that are more likely to exhibit multiple objects. It highlights the ongoing need for further advancements in reducing reliance on latent variables. To obtain a robust score and avoid the possible impact of the high variability of the seed’s success rate, we define a minimum of images to generate per prompt. We establish it is adequate to generate 32 images per prompt to obtain a robust score. The corresponding experiment is reported in the Supp. Mat. E.
@@ 4.3 Catastrophic Neglect
When examining the occurrence of different objects, similar trends can be observed as in the preliminary experiment (Tab. 2). Specifically, the initial objects in the template tend to appear more frequently than objects inserted subsequently. The results for the template containing four objects are presented in Fig. 5 (results for two and three objects in Supp. Mat. D). This reinforces the observation that the concept that is expressed earlier in the prompt has more chances to appear in the final image.
For SD, conditioned by CLIP text-encoders, the decreasing trend in the occurrence of objects as their position becomes more distant in the prompt may be partly due to the auto-regressive nature of the encoder. During self-attention, tokens only receive context from the elements to their left (beginning of the prompt). Tokens are thus devoid of the contexts of subsequent objects, while these last carry the context of the earlier words. Hence, the U-net model receives a more ambiguous signal from distant tokens.
The explanation is less clear with the T5 encoder, where all words have access to each other during self-attention. Used in IF, one can see that the third and fourth objects have the same chance to appear, but it is significantly less than the object in the second position, itself below the first one expressed in the prompt. We hypothesize that during cross-attention, the models learned to give more importance to tokens with earlier positions due to the training data, which typically places important elements related to the image at the beginning of the caption. Asserting this explanation would nevertheless require significant work to analyze the original training dataset used to pre-train the six generative models, which is out of the scope of this paper.
@@
Figure 5: The proportion of occurrences of each object, based on its position in the prompt (template with 4 objects).
@@
Finally, we investigated the impact of semantic relationships between objects within the prompt on the ability of the model to generate both of them. Considering a template with two objects, we hypothesized that the T2I model would fail more often to represent in the same image two objects that are semantically linked. We considered 28 COCO labels from three macro-classes (vehicles, animals, and foods) and generated images using a template with 2 objects. From the resulting TIAM scores, we derived a dissimilarity metric between all objects (see Supp. Mat. H) and projected all the labels with Multidimensional Scaling (MDS). The resulting projections (Fig. 6 for SD 1.4) can be interpreted such that the closer two labels are, the more challenging it becomes for the model to represent them together. Across all models, we observe labels from the same class being clustered together, particularly for animals and vehicles. This suggests the presence of semantic proximity. We measured the semantic distance between both named objects with various methods (such as Wu-Palmer, Cosine similarity of CLIP text embeddings, or even attention’s key-value representation) for SD 1.4 and SD 2 but the correlation with the TIAM score was only slightly negative. This indicates that semantic linking has probably either a small or indirect link to the alignment performance, but in any case, further research is needed to clarify this point.
@@ 4.4 Attribute binding
We start by computing TIAM regardless of the correctness of the attribute and report the results in Tab. 3. Without attributes (✗) the scores are close to those of Tab. 1 and adding attributes (✓) has a limited impact when one object only is requested. However, it significantly impacts the performance when two objects are present in the prompt.
@@
Figure 8: TIAM per color and object.
@@
When TIAM is estimated by taking into account both the object and color ground truth, we observe that the models fail to assign the appropriate colors to the objects, even when these objects are present in the image (Fig. 7). For instance, in the case of SD 1.4, objects are detected correctly in 95% of single-object prompts, but only approximately 45% of them have the requested colors.
In Fig. 8, we report the attribution scores per color and object to analyze deeper the attributing ability of the models. We see that models fail to generalize the binding to objects when it is uncommon to observe them in that particular color. As expected, it is indeed easier to assign colors to cars and refrigerators compared to animals. The models have likely been trained on photos of cars and refrigerators in various colors during training as it is more common to come across a green car than a green giraffe.
Following the same approach as in Section 4.3, we also analyze the results concerning the position $i$ of attribute $a_{i}$ and object $o_{i}$ in the prompt involving 2 objects. In that case, the models succeed by a large margin to generate and bind the first object. However, knowing that the first object is more often generated, we compute a binding success rate, which is the score of correctly attributed objects among the correctly generated objects (Fig. 9). The $o_{2}$ objects, however, continue to be less attributed. This reinforces the finding that the first object in the prompt has a greater influence on the final generation.
@@
In the Supp. Mat. I, we report the results of the binding success rate differentiated by colors for attributes in the first position and attributes in the second position. We observed that the models face greater difficulty in assigning green and blue colors when two objects are involved (parallel with a single label case). It is worth noting that IF performs better than other models.
4.5 Comparison to human, CLIP, BLIP
We randomly chose 32 prompt-image pairs and asked 57 humans to assess the content alignment. Their reliability of agreement was 0.73 in terms of Fleiss’kappa [12], which can be considered as “Substantial” [26]. Half of the prompts had two objects and the 16 others had one colored object. The Pearson correlation of humans and TIAM was 0.82 ($p<10^{-8}$). We compared TIAM to two other automatic methods based on CLIP [35] and BLIP [29], which had respectively a correlation of 0.47 ($p<10^{-2}$) and 0.67 ($p<10^{-4}$) with humans. As detailed in the Supp. Mat. K, TIAM had a better alignment with humans both for images with two objects and with one colored object.
@@ 4.6 Toward noise mining?
To highlight the importance of noise performance, we selected seeds based on their TIAM score. In Fig. 10, we present qualitative results using some of the worst and best seeds found with our approach, as well as prompts with objects and attributes that were not considered in our study.
Both with SD 2 and IF, the two objects are better represented in the image resulting from the “good” seed, showing that our method has the ability to find seeds that generalize to objects out of the domain of objects used to determine the best seeds. We conducted the same seed selection process for both our worst and best seeds using colors in the prompt (Fig. 11). Once again the images resulting from the best seeds we identified with TIAM better reflect the prompt. With the “bad” seed, SD 2 suffers both from attribute leaking or binding (the “blue moon” is yellow, and the “red lion” is blue), and for both models, the images resulting from the “good” seed are more aligned. This emphasizes the dependency on the noise present at the beginning of the prompt, which remains a prominent factor of performance.
@@
5 Conclusion, Limits and Perspective
We proposed a new metric to automatically quantify the performance of T2I models in terms of prompt-image alignment. Contrary to previous research efforts, it is based on prompt templates, that allow a finer analysis with regard to the syntax of the prompt. Hence, we showed that the alignment performance of most T2I models drops significantly with the number of objects specified in the prompt and that effect is even more critical for color attribution. Extending TIAM with more objects and attributes would result in the need to generate an exponential number of prompts, thus becoming cumbersome in practice. However, we show in Section L.1 of the Supp. Mat. that TIAM can be reliably estimated from a set of $\approx 300$ prompts. Our metric also allows us to study the influence of the input seed at inference. We showed that there exist some seeds that systematically result in better output images than others and that it generalizes to objects out of the set used to determine them. It draws possible future research toward the mining of such “good seeds”, similarly to some studies for text models [10], as a complementary activity to prompt engineering to optimize the outputs of T2I models. For a comprehensive analysis of T2I performance, TIAM should be combined with other metrics reflecting other aspects than prompt-image alignment, such as [42, 17, 7, 41, 51].